OpenStudy (anonymous):

What is the focus of the parabola given by the equation x = 3(y – 1)^2 + 2?

OpenStudy (amistre64):

i believe itll be related to x = 2y

OpenStudy (amistre64):

maybe x = 2(y-1) +2 to be centered right tho

OpenStudy (amistre64):

i forget what the geometric equations for these are ...

OpenStudy (amistre64):

3(y – 1)^2 + 2 = 2(y – 1) + 2 3(y – 1)^2 = 2(y – 1) 3(y – 1) = 2 y = 2/3 + 1 = 5/3, but id need to wolf it to be sure

OpenStudy (jdoe0001):

so it has a squared "y" component, thus the "focus form" of it will be \(\bf (y-k)^2=4p(x-h)\) (h, k) = vertex point p = distance from the vertex to the focus p < 0, opens to the left p > 0, opens to the right

OpenStudy (amistre64):

if im right ... and i seldom am: (2,5/3) maybe?

OpenStudy (amistre64):

yeah, im wrong .... its just turned all backwards is all

OpenStudy (jdoe0001):

well, came out as \(\bf x = 3(y-1)^2+2 \implies \cfrac{1}{3}(x-2) = (y-1)^2\\ 4p = \cfrac{1}{3} \implies p = \cfrac{1}{12}\)

OpenStudy (anonymous):

of the answer choices there is (2, 1 1/12) and (2 1/12, 1) those are the only that have 1/12

OpenStudy (jdoe0001):

so the vertex is at (2, 1) so the focus will be at \(\bf \left(2+ \cfrac{1}{12}, 1\right)\)

OpenStudy (amistre64):